# MemEvo: Memory-Evolving Incremental Multi-view Clustering

### v1

@@ Introduction
Table 1 shows the performance of the proposed method on six datasets with twelve comparison methods. 1) Despite being designed for incremental learning, the present method outperforms most traditional non-incremental methods. Notably, the proposed approach yields performance gains of 23.06% and 22.2% in ACC over the second-best method on ProteinFold and GRAZ02 datasets, respectively. 2) Our method demonstrates improvements over existing incremental methods (CMVC, LAIMVC, FDMVC, FCMVC). These improvements can be attributed to the balance of VAM and KCM, which addresses the plasticity-stability dilemma in incremental learning, thereby enabling efficient knowledge accumulation in incremental tasks. To enhance the validity of the results, we have provided code in the supplementary materials to facilitate reproduction.
Parametric Analysis
Sensitivity analysis of $\alpha$ and $\beta$.
To analyze the sensitivity of parameters $\alpha$ and $\beta$, we utilize the GRAZ02 and handwritten datasets as examples. With other parameters fixed, we investigate how variations within the interval $\{10^{-3},10^{-2},10^{-1},10^{0},10^{1}\}$ affect the accuracy (ACC) of our method. As shown in Figure 3(a), our method demonstrates low sensitivity to parameter variations when $\alpha$ and $\beta$ are within $\alpha,\beta\in\{10^{-3},...,10^{-1}\}$.
(a) Sensitivity analysis of $\alpha$ and $\beta$ on two datasets.
Sensitivity analysis of $m$.
To investigate the effect of latent space dimension, we conducted experiments by varying the parameter $m$ in the range of $[5,10,20,30,50]$. From Figure LABEL:m, we can observe that MemEvo exhibits significant robustness to dimensional changes and maintains stable performance across a wide range.
(b) The impact of $m$ and $\lambda$ on GRAZ02 dataset.
Sensitivity analysis of forgetting rate $\lambda$.
To investigate the impact of the forgetting rate $\lambda$, we conducted experiments on the GRAZ02 dataset, comparing different forgetting rate parameters within the range [0, 1, 1.5, 2]. As shown in Figure LABEL:lambda, when $\lambda=0$, the model degenerates to historical representation averaging, yielding the worst performance due to excessive smoothing of historical representation. Better performance occurs when $\lambda\in[1,2]$, indicating that controlled forgetting effectively filters redundant historical information while preserving useful knowledge.
Running time analysis.
@@
To evaluate the computational efficiency of our method, we compare MemEvo with other incremental methods in Table 2 in terms of runtime on four datasets. Although LAIMVC performs excellently in terms of speed, MemEvo achieves a balance between performance and efficiency in all scenarios. Notably, unlike CMVC and FCMVC, which either fail to complete (OM) or require significantly more time on large-scale datasets, our method successfully handles challenging cases such as Caltech101-all and MNIST.
Times(s) GRAZ02 handwritten Caltech101-all MNIST
CMVC 0.44 0.63 35.56 53.36
LAIMVC 0.07 0.01 0.28 1.79
FDMVC 0.93 0.78 4.67 12.05
@@
Ours 0.22 0.17 2.00 18.45
Table 2: Running time comparison on four datasets.
Representation Visualization Analysis
To show the quality of the representations learned, we visualized the GRAZ02 dataset in Figure 5(a). As can be seen from the figure, our method performs better in terms of intra-class cohesion and inter-class dispersion. This is because our collaborative strategy allows the model to retain the historical core knowledge while learning new information, which is crucial for generating high-quality representations.
(a) T-SNE visualizations of learned representation.
Ablation experiments
To evaluate the contribution of each module to model performance, we conducted ablation experiments on four datasets in Table 3. Our experiments reveal the contribution of each module: using only reconstruction loss yields poor performance, confirming its limited discriminative power. Adding VAM brings modest gains, demonstrating its role in view alignment. KCM drives major improvements, proving its long-term knowledge consolidation capability. Furthermore, CFM also contributes to improvement, suggesting that selectively discarding outdated knowledge can benefit incremental tasks. The collaborative effect of the three modules enables the model to achieve better clustering performance.
Methods Flower17 GRAZ02 Caltech101-all MNIST
$\mathcal{L}_{recon}$ 40.00 38.25 17.06 35.78
$\mathcal{L}_{recon}$+VAM 42.13 38.75 17.29 37.74
$\mathcal{L}_{recon}$+KCM 87.06 77.37 62.03 60.55
Ours (w/o CFM) 91.47 85.50 63.81 88.10
Ours 94.85 92.28 67.10 99.87
@@
Impact of view increments
Figure 7(a) shows the trend in model performance as the number of views increases, with the single-view training results used as a baseline for comparison. Theoretically, as new views are introduced, clustering performance should exhibit a monotonically increasing trend. Experimental results indicate that MemEvo consistently shows gradual improvement as the number of views increases. Notably, our incremental method shows a significant advantage over single-view learning. This phenomenon indicates that our model not only effectively transfers knowledge from existing views but also continuously integrates information from new views.
(a) Impact of view number on two datasets.
Conclusion
This paper proposes an incremental multi-view clustering framework based on memory evolution. Our core contribution lies in designing a brain science-inspired component architecture that integrates cognitive forgetting, view alignment, and knowledge consolidation memory modules to achieve a balance between model plasticity and stability. Compared with current state-of-the-art incremental clustering methods, our model demonstrates significant advantages across multiple evaluation metrics, validating its strong potential for dynamic multi-view clustering tasks. Although the current model performs well in existing incremental scenarios, our model has not explored more challenging dynamic learning scenarios, such as class-incremental learning. In future research, we will develop an automatic class discovery model based on uncertainty estimation to eliminate dependence on the number of predefined classes.

### v2

@@ 5.1 Experimental Results and Analysis
Table 2 shows the performance of the proposed method on six datasets with twelve comparison methods. 1) Despite being designed for incremental learning, the proposed method outperforms most traditional non-incremental methods. Notably, the proposed approach achieves performance gains of 26.16% and 9.94% in ACC over the second-best method on ProteinFold and Handwritten datasets, respectively. 2) Our method demonstrates improvements over existing incremental methods. These improvements can be attributed to the balance of SPD, thereby enabling efficient knowledge accumulation in incremental tasks.
5.2 Representation Visualization Analysis
To evaluate the quality of the learned representations, we visualize the GRAZ02 dataset. As observed in Figure 2, our method performs better in terms of intra-class cohesion and inter-class dispersion. This is because our collaborative strategy allows the model to retain the historical core knowledge while learning new information, which is crucial for generating high-quality representations.
(a) IWTSN (ACC = 73.31%)
(b) Ours (ACC = 92.28%)
Figure 2: T-SNE visualizations of learned representation.
5.3 Running Time analysis.
@@ 5.2 Representation Visualization Analysis
To evaluate the computational efficiency of our method, we compare EMIMC with other incremental methods in Table 3 on four datasets. Although LAIMVC excels in speed, EMIMC achieves a trade-off between performance and efficiency. Notably, unlike FCMVC, ASIA and CDIMVC, which fail to complete on large-scale datasets, our method successfully handles challenging cases such as Caltech101-all and FMNIST.
Table 3: Running time (s) comparison on four datasets.
Methods GRAZ02 Handwritten Caltech101-all FMNIST
CMVC 0.44 0.63 35.56 OM
LAIMVC 0.07 0.01 0.28 2.26
FDMVC 0.93 0.78 4.67 17.99
@@
ASIA 0.69 2.99 748.76 OM
CDIMVC 5.81 5.10 490.21 OM
Ours 0.22 0.17 2.00 14.30
5.4 Incremental View Analysis.
Impact of view increments. Figure 3 shows the trend in model performance as the number of views increases, with the single-view results used as a baseline for comparison. As new views are introduced, performance is expected to exhibit a monotonically increasing trend. Experimental results indicate that EMIMC consistently shows gradual improvement as the number of views increases. Notably, our method shows a significant advantage over single-view learning. This phenomenon indicates that our model not only effectively transfers knowledge from existing views but also continuously integrates information from new views.
(a) Flower17
(b) GRAZ02
Figure 3: Impact of view number on two datasets.
Impact of view Order. As indicated in Figure 3, significant performance variations exist among individual views. To investigate the impact of view input sequence on performance, we conducted experiments on the Flower17 and GRAZ02 datasets. Given the large number of available views, we evaluated three representative ordering scenarios: Max Order (sorted from best to worst individual performance), Original Order, and Min Order (sorted from worst to best). As shown in Figure 4, the results indicate that different ordering strategies have a minor impact on performance and consistently outperform the comparison method. This shows the robustness of our method to view order.
(a) Impact of View Order
(b) Convergence (FMNIST)
Figure 4: The impact of view order and Convergence curve.
5.5 Convergence Analysis
To demonstrate convergence, we plot the objective function values against the number of iterations. As observed in Figure 4, as the number of iterations increases, the experimental results on the two datasets indicate that the objective function value of the proposed algorithm shows a monotonically decreasing trend as iterations increase and finally converges to a stable point. Thus, the convergence of our method is guaranteed. Due to space constraints, the theoretical proof is shown in the supplementary materials.
5.6 Ablation Study
To evaluate the contribution of each module to model performance, we conducted ablation experiments on four datasets in Table 4. Our experiments demonstrate the contribution of each module: using only reconstruction loss yields poor performance, confirming its limited discriminative capability. Adding RAM brings modest gains, demonstrating its role in view alignment. KCM drives major improvements, proving its long-term knowledge consolidation capability. Furthermore, CFM also contributes to improvement, suggesting that selectively discarding outdated knowledge can benefit incremental tasks. The collaborative effect of the three modules enables the model to achieve better performance.
@@ 5.6 Ablation Study
Methods Flower17 GRAZ02 Caltech101-all FMNIST
$\mathcal{L}_{recon}$ 40.00 38.25 17.06 10.54
$\mathcal{L}_{recon}$+RAM 42.13 38.75 17.29 10.61
$\mathcal{L}_{recon}$+KCM 87.06 77.37 59.22 32.13
Ours (w/o CFM) 90.74 83.20 59.28 33.97
Ours 93.37 92.28 64.46 82.86
5.7 Parametric Analysis
Sensitivity analysis of $\alpha$ and $\beta$. To analyze the sensitivity of parameters $\alpha$ and $\beta$, we use the GRAZ02 and Handwritten datasets as representative datasets. With other parameters fixed, we investigate how variations within the interval $\{10^{-4},10^{-3},10^{-2},10^{-1},10^{0},10^{1}\}$ affect the accuracy (ACC) of our method. As shown in Figure 5, our method demonstrates low sensitivity to parameter variations when $\alpha$ and $\beta$ are within $\alpha,\beta\in[10^{-3},10^{-2}]$.
(a) GRAZ02
(b) Handwritten
Figure 5: Sensitivity analysis of $\alpha$ and $\beta$ on two datasets.
Sensitivity analysis of $m$. To investigate the impact of the latent space dimension, we conducted experiments by varying the parameter $m$ in the range of $[5,10,20,30,50]$. From Figure 6, we can observe that EMIMC exhibits significant robustness to dimensional changes and maintains stable performance across a wide range. In our experiment, we set $m=50$ for all datasets.
(a) Latent space dimension $m$
(b) Forgetting rate $\lambda$
Figure 6: The impact of $m$ and $\lambda$ on six datasets.
Sensitivity analysis of forgetting rate $\lambda$. To investigate the impact of the forgetting rate $\lambda$, we conducted experiments on the GRAZ02 dataset, comparing different forgetting rate parameters within the range [0, 0.5, 1, 1.5, 2]. As shown in Figure 6, when $\lambda=0$, the model degenerates to historical representation averaging, yielding the poorest performance due to excessive smoothing of historical representations. Better performance occurs when $\lambda\in[1,1.5]$, indicating that controlled forgetting filters redundant historical information while preserving useful knowledge. In our experiment, we just set $\lambda=1$ for all datasets.
6 Conclusion
This paper proposes an incremental multi-view clustering framework based on memory evolution. Our core contribution lies in designing a brain science-inspired component architecture to achieve a balance between plasticity and stability. Compared with current methods, our model demonstrates significant advantages across multiple evaluation metrics, validating its potential for incremental tasks. Although EMIMC performs well in existing incremental scenarios, our model has not explored more challenging scenarios, such as class-incremental learning. In future research, we will develop a class discovery model to eliminate dependence on the number of predefined classes.